Sum-of-Years' Digits, Declining Balance

Cost of Equipment: $750,000

Estimated Salvage Value: $70,000

Estimated Life (yrs): 16

Sold after this many years: 10

Selling price of equipment: $180,000

---------------------​​​​​​​

Please fill out the blanks bellow,

A.

Depreciable Cost
Sum of years digits (life)
Sum of years digits (yrs remaining)
Declining balance rate

B.

Depreciation Expense SYD Method DB Method
Year 2
Year of sale
At time of sale SYD Method DB Method
Accumulated Depreciation
Book Value

C.

ENTRY (SYD Method):
DR Account CR Account DR Amount CR Amount
Answer #1

A. Depreciable Cost= $750000-$70000= 680000

Sum of years digit (Life)= 1+2+3+4+5+6+7+8+9+10+11+12+13+14+15+16=136

Sum of years digit (Yrs remaining)= 11+12+13+14+15+16= 81

Declining balance rate= 100/16= 6.25%

B.

Depreciation exp

SYD method

DB Method

Year 2

75000 (750000-70000)x16/136

43945 (703125x6.25%)

Note: 750000-6.25%=703125

year of sale (i.e. for year 11)

30000 (750000-70000)x6/136

24584 (393345x6.25%)

At the time of sale

SYD method

DB method

Accumulated Depreciation

575000

356655

Book value

175000

393345

C. Entry under SYD Method:

Year 1 = Depreciation A/c. ---Dr. 80000

to Accumulated Depreciation / fixed assets A/c. 80000

Like the above entry ; each year depreciation entry will be done.

Depreciation schedule for the assets(for 10 years):

Year

Depreciation Base

Remaining Life of machine

Depreciation Fraction

Depreciation Expense

1

680000

16

16/136

80000

2

680000

15

15/136

75000

3

680000

14

14/136

70000

4

680000

13

13/136

65000

5

680000

12

12/136

60000

6

680000

11

11/136

55000

7

680000

10

10/136

50000

8

680000

9

9/136

45000

9

680000

8

8/136

40000

10

680000

7

7/136

35000